A bomb of mass 1 kg, thrown vertically upwards with a speed of u=100m/sec, explodes into two parts after t=5 secs. A fragment of mass m1= 400 g moves downwards with a speed v1= 25 m/sec, then speed v2 and direction of another mass m2 will be-
a) 40 m/s downwards  b) 40m/s upwards  c) 60m/s upwards  d) 100m/s upwards

Grade:11

1 Answers

Khimraj
5 years ago
velocity of bomb after 5 sec = 100 – 10*5 = 50 m/s
then applying conservation of momentum and taing upward direction as positive
1*50 = (2/5)*(-25) + (3/5)*x
then x = 100m/s upwards.
Hope it clears.